OpenStudy (anonymous):

The hypotenuse of a right triangle is 26 millimeters. One leg of the right triangle is 10 millimeters. What is the length of the other leg? A. 18 millimeters B. 20 millimeters C. 24 millimeters D. 28 millimeters

OpenStudy (anonymous):

Rearrange the pythagoream theorem to get the answer. a^2 + b^2 = c^2 c^2 - a^2 = b^2 c= 26 a = 10 26^2 - 10^2 = b^2 676 - 100 = b^2 576 = b^2 Square root both sides. b = 24 millimeters.

OpenStudy (anonymous):

thank youu veryyy much for helping (:

OpenStudy (anonymous):

No problem :)